kekeman:

Determine how many, what type, and find the roots for f(x) = x3 − 5x2 − 25x + 125.

surjithayer:

it is a cubic polynomial and has three roots. f(x)=0 ,gives \[x^3-5x^2-25x+125=0\] \[x^2(x-5)-25(x-5)=0\] \[(x-5)(x^2-25)=0\] \[(x-5)(x+5)(x-5)=0\] x=?

kekeman:

@surjithayer wrote:
it is a cubic polynomial and has three roots. f(x)=0 ,gives \[x^3-5x^2-25x+125=0\] \[x^2(x-5)-25(x-5)=0\] \[(x-5)(x^2-25)=0\] \[(x-5)(x+5)(x-5)=0\] x=?
x = 5, -5